4x-2 divided by 3 =-6
Answer by victoria.howard(6)

Solve using the multiplication principle.
4x-2/3 =-6
We need to isolate the variable x to one side of the equation.
To cancel out the -2/3 multiply by the inverse
4x-2/3*3/2 = -6/1*3/2
4x= -18/2
4x = -9
Divide both sides by 4
4x/4 = -9/4
x= -9/4